OpenStudy (anonymous):

A range of a set of numbers is 30. The highest number in the set is two times larger than the lowest number in the set. Which set of algebraic equations will solve for the highest and lowest numbers in the set?

OpenStudy (anonymous):

A on the grap is a little under 11 B is 9 C is 3 D is a little under 4

OpenStudy (liv1234):

x is going to equal the lowest number and y is going to be the highest number. the range is the highest minus the lowest number, so y-x=30. the highest is 2 times larger than the lowest so y=2x

OpenStudy (anonymous):

range = 30 = the difference between the largest number and the lowest number of the set of data If the lowest number in the set is x, then the larger one will be 2(30 - x). So we have: 2(30 - x) - x = 30 60 - 2x - x = 30 60 - 3x = 30 60 - 3x + 3x - 30 = 30 - 30 + 3x 30 = 3x 30/3 = 3x/3 10 = x the lowest number 2(30 - x) = 2(30 - 10) = 40 the largest number Y= high number, X= low number y-x=30 y=2x

OpenStudy (liv1234):

\[y = 2x\]

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

Your welcome.

OpenStudy (mrnood):

I don't think that is correct @Lovelarap @khorribrown

OpenStudy (mrnood):

the lowest number is x so the highest number is 2x the range is 30 so 2x-x =30 x=30 highest number = 2x = 60

OpenStudy (mrnood):

@Lovelarap 10 and 40 does not satisfy the y=2x equation I don't need to google it - I have worked it out

OpenStudy (mrnood):

@Lovelarap Your equations are correct: y-x=30 y=2x but the solution is 30, 60 not 10,40